Find the average speed of all runners having a speed of 2 m/s, 3m/s and 4 m/s respectively?

a. 2 m/s b. 3 m/s2 c. 3 m/s d. 4.5 m/s




1
Expert's answer
2021-03-29T09:06:42-0400
"v_{av}=\frac{2+3+4}{3}=3\frac{m}{s}"

Answer: c)